OpenStudy (anonymous):

can anyone help me out? I did/started the first three questions. I need the fourth. Answer first, then you can walk me through it c: Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A B C Price f(x) = 25(1.08)x x f(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day.

OpenStudy (anonymous):

1. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. f(x) = 26(1.3)x 2. Graph the price function for Stock D. Explain the key features of the graph with complete sentences. f(1) = 26(1.3)1 f(1) = 33.8? f(2) = 26(1.3)2 f(2) = 67.6 3. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). 1 + -6 2 + -6 3 + -6 4. Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. • f(x + 2) • f(–x) • f(x) + 3

OpenStudy (anonymous):

I need the 4th q

OpenStudy (anonymous):

@robtobey @lacrosseplayer22 @NaeNaeHoowaahh ?

OpenStudy (anonymous):

f(x+2) increases all x values by 2. f(-x) multiplies the x values by one. f(x)+3 is adding three to the line. At least i think this is right.